# Unforgettable Lessons from Forgettable Images: Intra-Class Memorability Matters in Computer Vision Tasks

### v1

@@ 4.1 Memorability Prediction
Memorability prediction can provide insights into human cognition and perception by helping us understand how and why certain images are more memorable than others [42, 19, 37]. Existing memorability prediction methods fail to disentangle the intrinsic and extrinsic factors that influence the memorability of an image. In contrast, our work focuses on predicting ICMscore based exclusively on intrinsic factors of images within the same category.
@@
Table 1: Performance comparison of ICMscore prediction among different AI models. The Spearman Rank Correlation [54] quantifies the alignment between predicted and ground truth ICMscores, where higher correlations denote more accurate prediction performance. The number of trainable parameters, the total number of model parameters, and the number of FLOPs (floating-point-operations) are shown in Row 2, 3, and 4 respectively. Numbers in the brackets on Row 1 are standard errors.
Specifically, we use a ResNet [20], pre-trained on ImageNet [11], to extract feature representations for an input image and then attach two fully connected layers to predict the memorability score. We train our ICMscore predictor on the ICMD dataset with an L2 regression loss, where the ICMscore of the input image serves as the ground truth. To validate the effectiveness of our predictor, we compare it against two baseline models. Baseline 1 employs the same network structure as our model but is trained from scratch without the pre-trained feature extractor. For Baseline 2, inspired by previous work [17], we use a Vision Transformer (ViT) [2] trained from scratch. We perform 5-fold cross-validation to ensure the statistical robustness of the results.
Following the established practices in memorability prediction [24, 17, 26], we assess the models’ performance by measuring the Spearman rank correlation [54] between the predicted and ground truth ICMscores. In Tab. 1, our model achieves the highest correlation (0.62 ± 0.06), demonstrating that ICMscores can be effectively predicted using features extracted from pre-trained image recognition models. These results indicate that features valuable for object recognition are also informative for predicting memorability. This finding aligns with our observation of the positive correlation between feature distance from the categorical centroids and ICMscores (Sec 3.3).
We detailed the number of trainable and total model parameters for our predictor and the two baselines in Row 2 and 3 of Tab. 1. Despite achieving accurate memorability predictions, our predictor has the fewest trainable parameters (0.39M) and the smallest model size (21.68M) among all the predictors, suggesting that larger models do not always yield better performance. Furthermore, our predictor requires the least number of FLOPs (3.68G), demonstrating its high computational efficiency.
To further explore what part of the image contributes to high memorability, we use Grad-CAM [46] to visualize the importance maps for memorability, shown in Fig. 4. For HM images, our predictor focuses on distinctive regions. For instance, in a teapot image with an ICMscore of 0.78, our predictor emphasizes the unusual face texture of the teapot, while in a flower image with an ICMscore of 0.75, the model highlights the vibrant yellow flower in contrast to the smaller pale white flowers in the background. This suggests that exaggerated deviations from typical visual features make an image more memorable.
Figure 4: Visualization of the importance maps on two example images with high memorability predicted by our model based on Grad-CAM [46]. For HM images, our ICMscore predictor highlights informative regions that contribute to high ICMscores. Qualitatively, these highlighted regions correspond to salient visual features that enhance memorability. The activation values from Grad-CAM are indicated by the color bar on the right.
@@ 4.2 Image Recognition
We explore the role of memorability in image recognition by training models on image subsets with varying ICMscores. Images within each object class are sorted by their ICMscores and divided into two subsets: HM (High Memorability) with the top 250 images and LM (Low Memorability) with the bottom 250. Additionally, we create a Mixed-Memorability (MM) subset of 250 images randomly sampled from the ICMD, containing both HM and LM images.
@@
We employed ResNet [20] as the backbone for object recognition, training it from scratch on each of the three curated subsets using the Adam optimizer [27]. The initial learning rate was set to $1\times 10^{-3}$ and reduced by 50% every two epochs over a total of 10 epochs following a linear decay schedule. The resulting models were labeled as HM-model, MM-model, and LM-model based on their training subsets.
@@
We used the Cross-Entropy loss to train all three models for the 5-way classification task. Test images for object recognition were sourced from the original ImageNet [11] test set. However, as these test images lack ground truth ICMscores, we could not evaluate the top-1 accuracy of the models on subsets of test images stratified by ICMscores. Figure 5 presents the average top-1 accuracies of the three models on all the test images for all 5 object classes over 5 random seeds.
Figure 5: Performance comparison in Top-1 Accuracy between image classification models trained on HM, MM, and LM images. From the top to the bottom, Top-1 Accuracy is reported for models trained on LM images (blue), HM images (green) and MM images (yellow). The chance is 20% for the 5-way classification (dash lines produced by the random model in red). The error bar is the standard error calculated over 5 independent runs with different random seeds. One-way ANOVA across the results of all models yields p $<0.05$, indicating statistical significance in accuracy differences among green, yellow, and blue bars.
@@
Our results in Figure 5 indicate that the LM-model consistently outperforms the HM-model, achieving an average top-1 accuracy of $0.83$ compared to $0.77$ across all 5 object classes. Statistical analysis using one-way ANOVA revealed significant differences in top-1 accuracy between the two models ($p<0.05$). Surprisingly, the LM-model also surpasses the MM-model. These findings indicate that image memorability impacts image recognition performance, with LM images improving accuracy and HM images impairing it. We also provide the confusion matrices for the classification performances of all three models in Sec. S4.
@@
In real-world applications, evaluating the memorability of all images through human behavioral studies is often impractical before training recognition models with images of varying memorability. Therefore, a computational model for predicting an image’s memorability becomes invaluable. We leveraged the ICMscore predictor introduced in Sec. 4.1 to estimate the ICMscore for all images and then used these predictions as criteria to conduct the same experiment for image recognition. The results are shown in Fig. S6 (a). Consistent with findings based on ICMscores derived from human behavioral experiments, recognition models trained exclusively on low-memorability (LM) images categorized by predicted ICMscores achieved higher top-1 accuracies than the MM-model from our ICMD dataset.
@@
For comparison, we also conducted the same image recognition experiment using memorability definitions from [7, 26, 13, 15, 30] to divide the dataset into two subsets and utilizing the predictor from [26]. The results, shown in Fig. S6 (b), indicate that the predictor proposed by [26] does not enhance image recognition accuracy. Additionally, we observed no significant difference in top-1 accuracy between models trained on the HM and LM images as predicted by [26].
@@ 4.3 Continual Learning
Continual learning, particularly class-incremental learning, is essential for AI models to adapt to new classes over time without experiencing catastrophic forgetting of previously learned classes [39]. One type of continual learning methods is based on memory replay [3, 50, 47], which involves storing a subset of previously encountered examples from old classes in a memory buffer and replaying them during subsequent tasks. Given that some images are easier to remember while others are more easily forgotten, we conjecture that image memorability may play a key role in improving the continual learning performance of AI models. Here, we investigate how memorability can be leveraged to select exemplar images for replay, potentially enhancing the model’s ability to retain knowledge across tasks.
Figure 6: Continual learning performance of AI models using HM, MM, and LM images for replays. From left to right, the performance of continual learning models using HM (green), MM (yellow), and LM (blue) images for replays is presented in Acc (left), BWT (middle), and Forgetting (right). See Sec. 4.3 for the introduction to the evaluation metrics and the continual learning models. The error bars are the standard errors calculated over 5 independent runs with different random seeds. A two-tailed t-test analysis is performed to compare the performance in all three evaluation metrics between HM-models and LM-models over 5 runs (p-values $<$ 0.05) and between MM-models and LM-models over 5 runs (p-values $<$ 0.05).
@@
Accuracy at the last task (Acc) [47] is defined as the accuracy of the model trained up to the last task $T$ and tested on all the previous tasks. The higher the Acc, the better the model performs in continual learning tasks.
Backward Transfer (BWT)[33, 12, 47] quantifies the model’s knowledge retention ability by measuring performance drop on earlier tasks after learning new ones:
@@
Forgetting (F) [47] captures the accuracy drop on the test set of the first task for the model trained at the final task. It is defined as $F=A_{1,1}-A_{T,1}$. The lower the F, the better the model performs in continual learning tasks.
Results in Fig. 6 show that models using LM images for replay consistently outperform those using HM and MM images in Acc and BWT ($p<0.05$). Additionally, models using LM images exhibit reduced forgetting, indicated by lower F, though this difference is not statistically significant. We also plotted Acc and F as functions of task numbers for all three models in Fig. 7. The results indicate that the model replayed with LM images consistently exhibits reduced forgetting and achieves higher accuracy across all tasks. Together with our feature analysis in Sec. 3.3, our results in continual learning experiments align well with the previous research [43, 25] that replaying samples where their features are closer to the categorical centroids yields better continual learning performance.
@@
Similar to the image recognition experiments in Sec. 4.2, we categorized images into HM, MM, and LM using both our ICMscore predictor and the predictor from [26]. From Fig. S6, there is no performance difference in replay methods using LM and HM images predicted by [26]. In contrast, our ICMscore predictor remarkably reveals significant differences, with continual learning models replayed with LM images outperforming those with HM images or models using replay images predicted by [26].
@@ 4.4 Memorability-controlled Image Editing
Editing images with controlled memorability offers valuable applications across various research fields, such as designing advertisements for commercial products and improving students’ knowledge retention ability in education [16, 14, 36]. Next, we present a proof-of-concept demonstration of using a subset of teapot images in our ICMD dataset to train a state-of-the-art image diffusion model capable of modifying the memorability of an original teapot image.
To adjust a teapot image’s memorability while preserving its semantic content, we trained a high-memorability teapot generator (HM generator) using a subset of teapot images with high memorability (HM images) calculated from human behavioral experiments. First, for each HM image, we used SAM [28] to segment and mask the teapot area with white pixels. These masked images were paired with their corresponding original images to train image diffusion models. Specifically, we fine-tuned the pre-trained InstructPix2Pix model (Ip2p) [6] with LoRA [22]. The HM generator uses “highly memorable teapot” as the text prompt and the masked teapot image as input for image editing. The masked area guides the HM generator to focus on the specified region. During fine-tuning, the generator produces highly memorable teapot images and compares them against the ground truth HM images to improve its performance.
@@
Figure 8: Proof-of-conept demonstration of memorability-controlled image editing. (a) During training, we fine-tune the image diffusion model, Ip2p  [6], using LoRA [22] with pairs of masked images and their corresponding ground truth images. (b) During inference, SAM [28] segments and masks the teapot area of the source image, and our fine-tuned model takes the masked image as input to edit teapot images with varying memorability. (c) The high-memorability image generator increases the ICMscore of low-memorability (LM) teapots by adding striking textures and enhancing brightness. Conversely, the low-memorability image generator reduces the ICMscore by blurring distinctive features, making the teapots appear more ordinary with plain textures. The numbers beneath the images represent ICMscores predicted by our memorability predictor as introduced in Sec. 4.1.
5 Conclusion and Future Works
@@
We introduce intra-class memorability, defined as the ability of individual instances within the same image class to be remembered by humans. To investigate why certain instances are more memorable than others, we contributed the ICMD dataset, comprising 2,500 images across 5 object classes. Memorability for these images is quantified using the ICMscore metric, derived from human behavioral experiments in continuous object recognition tasks.
To showcase the utility of ICMD, we train and evaluate AI models on subsets of the dataset, categorized by ICMscore, across four downstream computer vision tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing. Our results reveal that intra-class memorability significantly impacts feature representation learning, with low-memorability images notably improving performance in image recognition and continual learning tasks.
We identify several promising directions for future research in intra-class memorability. While our work provides proof-of-concept demonstrations of memorability’s role in various computer vision tasks, more rigorous experiments could further explore its potential. For instance, in memorability-controlled image editing, more precise controls on memorability could be imposed at finer-grained levels during the image editing process. Additionally, while our study primarily focused on supervised training, future research could investigate the impact of intra-class memorability within self-supervised learning frameworks.
Supplementary Material
S1 Difference in Human Behavior Experiments from the Existing Literature

### v2

@@ 4.1 Memorability Prediction
Memorability prediction can provide insights into human cognition and perception by helping us understand how and why certain images are more memorable than others (Perera et al., 2019; Harada & Sakai, 2023; Needell & Bainbridge, 2022). Existing memorability prediction methods fail to disentangle the intrinsic and extrinsic factors that influence the memorability of an image. In contrast, our work focuses on predicting ICMscore based exclusively on intrinsic factors of images within the same category.
Specifically, we use a ResNet (He et al., 2015), pre-trained on ImageNet (Deng et al., 2009) for image recognition tasks, to extract feature representations for an input image and then attach two fully connected layers to predict the memorability score. We train our ICMscore predictor on the ICMD dataset with an L2 regression loss, where the ICMscore of the input image serves as the ground truth. To validate the effectiveness of our predictor, we compare it against two baseline models. Baseline 1 employs the same network structure as our model but is trained from scratch without the pre-trained feature extractor on ImageNet for image recognition tasks. For Baseline 2, inspired by previous work (Hagen & Espeseth, 2023; Constantin & Ionescu, 2021), we use a Vision Transformer (ViT) (Alexey, 2020) trained from scratch. We perform 5-fold cross-validation to ensure the statistical robustness of the results.
Following the practices in memorability prediction (Isola et al., 2014; Hagen & Espeseth, 2023; Khosla et al., 2015), we assess the models’ performance by measuring the Spearman rank correlation (Zar, 2005) between the predicted and ground truth ICMscores. In Tab. 1, our model achieves the highest correlation (0.62 ± 0.06), demonstrating that ICMscores can be effectively predicted using features extracted from pre-trained image recognition models.
To ensure the 0.62 correlation from our predictor is not due to model overfitting, we split human participants into two independent groups and computed the Spearman correlation of their ICMscores on the same target images. Across 10 runs, the average correlation was $0.69\pm 0.05$, demonstrating high consistency in within-human memorability. While our predictor’s correlation is slightly lower than human consistency, it surpasses both baselines, suggesting that features crucial for object recognition also contribute to predicting memorability. This finding aligns with the observed positive correlation between feature distance from the categorical centroids and ICMscores (Sec 3.3).
Table 1: Performance comparison of ICMscore prediction among different AI models. The Spearman Rank Correlation (Zar, 2005) quantifies the alignment between predicted and ground truth ICMscores. The higher, the better. The number of trainable parameters, the total number of model parameters, and the number of FLOPs (floating-point-operations) are also presented. Numbers in the brackets are standard errors.
@@
We detailed the number of trainable and total model parameters for our predictor and the two baselines in Tab. 1. Despite achieving accurate memorability predictions, our predictor has the fewest trainable parameters and the smallest model size, suggesting that larger models do not always yield better performance. Furthermore, our predictor requires the least number of FLOPs, demonstrating its high computational efficiency.
To further explore what part of the image contributes to high memorability, we use Grad-CAM (Selvaraju et al., 2017) to visualize the importance maps for memorability, shown in Fig. 4. For HM images, our predictor focuses on distinctive regions. For instance, in a teapot image with an ICMscore of 0.78, our predictor emphasizes the unusual face texture of the teapot, while in a flower image with an ICMscore of 0.75, the model highlights the vibrant yellow flower in contrast to the smaller pale white flowers in the background. This suggests that exaggerated deviations from typical visual features make an image more memorable.
Figure 4: Visualization of the importance maps on two example images with high memorability highlighted by our ICMscore predictor based on Grad-CAM (Selvaraju et al., 2017). Grad-CAM visualizes key image regions by computing gradients of our ICMscore predictor against ground truth scores. For HM images, it highlights features that enhance memorability, with activation values shown in the color bar on the right.
@@ 4.2 Image Recognition
We explore the role of memorability in image recognition by training models on image subsets with varying ICMscores. Images within each object class are sorted by their ICMscores and divided into two subsets: HM (High Memorability) with the top 250 images and LM (Low Memorability) with the bottom 250. Besides, we create a Mixed-Memorability (MM) subset of 250 images randomly mixing HM and LM images from the ICMD.
@@
We employed ResNet (He et al., 2015) as the backbone with random weights, and trained it from scratch on each of the three curated subsets above using the Adam optimizer (Kingma & Ba, 2017). The initial learning rate was set to $1\times 10^{-3}$ and reduced by 50% every two epochs over a total of 10 epochs following a linear decay schedule. The resulting models were labeled as HM-model, MM-model, and LM-model based on their training subsets.
@@
We used the Cross-Entropy loss to train all three models for the 5-way classification task. Test images for object recognition were sourced from the original ImageNet (Deng et al., 2009) test set. However, as these test images lack ground truth ICMscores, we could not evaluate the top-1 accuracy of the models on subsets of test images stratified by ICMscores. Fig. 5 presents the average top-1 accuracies of the three models on all the test images for all 5 object classes over 5 random seeds.
@@
Our results in Fig. 5 indicate that the LM-model consistently outperforms the HM-model, achieving an average accuracy of $0.83$ compared to $0.77$ across all 5 object classes. Statistical analysis using one-way ANOVA revealed significant differences in top-1 accuracy between the two models ($p<0.05$). Surprisingly, the LM-model also surpasses the MM-model. These findings indicate that image memorability impacts image recognition, with LM images improving accuracy and HM images impairing it. We also provide the confusion matrices for the classification performances of all three models in Sec. A.4.
Figure 5: Performance comparison in Top-1 Accuracy between image classification models trained on HM, MM, and LM images. From the top to the bottom, Top-1 Accuracy is reported for models trained on LM images (blue), HM images (green) and MM images (yellow). The chance is 20% for the 5-way classification (red dash line). The error bar is the standard error calculated over 5 independent runs. One-way ANOVA across the results of all models yields p $<0.05$, indicating statistical significance in accuracy differences among green, yellow, and blue bars.
@@
In real-world applications, evaluating the memorability of all images through human behavioral studies is often impractical before training recognition models with images of varying memorability. Therefore, a computational model for predicting an image’s memorability becomes invaluable. We leveraged the ICMscore predictor (Sec. 4.1) to estimate the ICMscore for all images and then used these predictions as criteria to conduct the same experiment for image recognition. From Fig. A6 (a), consistent with findings above, recognition models trained exclusively on LM images categorized by predicted ICMscores achieved higher top-1 accuracies than the MM-model from our ICMD dataset. One possible reason is that easy images for object recognition often feature prototypical representations of object classes (Singh et al., 2023), which accelerate training and foster robust, generalizable learning. As shown in Sec 3.3, easy images tend to be nearer to categorical centroids, and hence, less memorable but more beneficial for learning.
@@
For comparison, we also conducted the same image recognition experiment using memorability definitions from (Bylinskii et al., 2015; Khosla et al., 2015; Dubey et al., 2015; Goetschalckx & Wagemans, 2019; Kramer et al., 2023) to divide the dataset into two subsets and utilizing the predictor from (Khosla et al., 2015). The results, shown in Fig. A6 (b), indicate that the predictor proposed by (Khosla et al., 2015) does not enhance image recognition accuracy. Additionally, we observed no significant difference in top-1 accuracy between models trained on the HM and LM images as predicted by (Khosla et al., 2015).
@@ 4.3 Continual Learning
Continual learning, particularly class-incremental learning, is essential for AI models to adapt to new classes over time without experiencing catastrophic forgetting of previously learned classes (Parisi et al., 2019). One type of continual learning methods is based on memory replay (Bagus & Gepperth, 2021; Tee & Zhang, 2023), which involves storing a subset of previously encountered examples from old classes in a memory buffer and replaying them during subsequent tasks. Given that some images are easier to remember while others are more easily forgotten, we conjecture that image memorability may play a key role in improving the continual learning performance of AI models. Here, we investigate how memorability can be leveraged to select exemplar images for replay, potentially enhancing the model’s ability to retain knowledge across tasks.
Figure 6: Continual learning performance of AI models using HM, MM, and LM images for replays. From left to right, the performance of continual learning models using HM (green), MM (yellow), and LM (blue) images for replays is presented in Acc (left), BWT (middle), and Forgetting (right). See Sec. 4.3 for the introduction to the evaluation metrics and the continual learning models. The error bars are the standard errors calculated over 5 independent runs with different random seeds. A two-tailed t-test analysis is performed to compare the performance in all three evaluation metrics between HM-and LM-models over 5 runs and between MM- and LM-models over 5 runs (p-values $<$ 0.05).
@@
Accuracy at the last task (Acc) (Shi et al., 2024) is defined as the accuracy of the model trained up to the last task $T$ and tested on all the previous tasks. The higher the Acc, the better the model performs in continual learning tasks.
Backward Transfer (BWT)(Lin et al., 2022; Díaz-Rodríguez et al., 2018; Shi et al., 2024) quantifies the model’s knowledge retention ability by measuring performance drop on earlier tasks after learning new ones:
@@
Forgetting (F) (Shi et al., 2024) captures the accuracy drop on the test set of the first task for the model trained at the final task. It is defined as $F=A_{1,1}-A_{T,1}$. The lower the F, the better the model performs in continual learning tasks.
Results in Fig. 6 show that models using LM images for replay consistently outperform those using HM and MM images in Acc and BWT ($p<0.05$). Additionally, models using LM images exhibit reduced forgetting, indicated by lower F, though this difference is not statistically significant. We also plotted Acc and F as functions of task numbers for all three models in Fig. 7. The results indicate that the model replayed with LM images consistently exhibits reduced forgetting and achieves higher accuracy across all tasks. Together with our feature analysis in Sec. 3.3, our continual learning results align well with the previous research (Rebuffi et al., 2017; Ke et al., 2024) that replaying samples where their features are closer to the categorical centroids yields better continual learning performance.
@@
Similar to the image recognition experiments in Sec. 4.2, we categorized images into HM, MM, and LM using both our ICMscore predictor and the predictor from (Khosla et al., 2015). From Fig. A6, there is no performance difference in replay methods using LM and HM images predicted by (Khosla et al., 2015). In contrast, our ICMscore predictor remarkably reveals significant differences, with continual learning models replayed with LM images outperforming those with HM images or models using replay images predicted by (Khosla et al., 2015).
@@ 4.4 Memorability-controlled Image Editing
Editing images with controlled memorability offers valuable applications across various research fields, such as designing advertisements for commercial products and improving students’ knowledge retention ability in education (Goetschalckx et al., 2019; Garcia-Retamero et al., 2012; Moran et al., 2019). Next, we present a proof-of-concept demonstration of using a subset of teapot images in our ICMD dataset to train a state-of-the-art image diffusion model capable of modifying the memorability of an original teapot image.
To adjust a teapot image’s memorability while preserving its semantic content, we trained a high-memorability teapot generator (HM generator) using a subset of teapot images with high memorability (HM images) calculated from human behavioral experiments. First, for each HM image, we used SAM (Kirillov et al., 2023) to segment and mask the teapot area with white pixels. These masked images were paired with their corresponding original images to train image diffusion models. Specifically, we fine-tuned the pre-trained InstructPix2Pix model (Ip2p) (Brooks et al., 2022) with LoRA (Hu et al., 2021). The HM generator uses “highly memorable teapot” as the text prompt and the masked teapot image as input for image editing. The masked area guides the HM generator to focus on the specified region. During fine-tuning, the generator produces highly memorable teapot images and compares them against the ground truth HM images to improve its performance.
@@
Figure 8: Proof-of-conept demonstration of memorability-controlled image editing. (a) During training, we fine-tune the image diffusion model, Ip2p  (Brooks et al., 2022), using LoRA (Hu et al., 2021) with pairs of masked images and their corresponding ground truth images. (b) During inference, SAM (Kirillov et al., 2023) segments and masks the teapot area of the source image, and our fine-tuned model takes the masked image as input to edit teapot images with varying memorability. (c) The high-memorability image generator increases the ICMscore of low-memorability (LM) teapots by adding striking textures and enhancing brightness. Conversely, the low-memorability image generator reduces the ICMscore by blurring distinctive features, making the teapots appear more ordinary with plain textures. The numbers beneath the images represent ICMscores predicted by our memorability predictor as introduced in Sec. 4.1.
5 Conclusion and Future Work
@@
We introduce intra-class memorability, defined as the ability of individual instances within the same image class to be remembered by humans. To investigate why certain instances are more memorable than others, we contributed the ICMD dataset, comprising 2,500 images across 5 object classes. Memorability for these images is quantified using the ICMscore metric, derived from human behavioral experiments in continuous object recognition tasks. To showcase the utility of ICMD, we train and evaluate AI models on subsets of the dataset, categorized by ICMscore, across four downstream computer vision tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing. Our results reveal that intra-class memorability significantly impacts feature representation learning, with low-memorability images notably improving performance in image recognition and continual learning tasks.
Building on our promising results, we outline several key directions for future research. While our study demonstrates memorability’s role in computer vision using five object classes, expanding ICMD to hundreds of ImageNet classes would provide broader insights. Furthermore, controlling image properties—such as standardizing backgrounds—could remove confounding factors and help isolate class-specific features critical for memorability. Although our findings highlight memorability’s impact on computer vision tasks, its underlying mechanisms remain an open question. Lastly, while we focused on supervised training, future work could explore intra-class memorability in self-supervised learning frameworks.
Impact Statement
We introduce the new problem of intra-class memorability and establish the first framework to systematically study and assess its features in class-specific naturalistic images. Notably, low-memorability images enhance object recognition and continual learning. Our work also enables generative AI to change image memorability, with potential applications in education and commercialization. However, it could be misused to manipulate the attention and memory of the general public, highlighting the need for ethical guidelines and safety regulations.
Appendix A Appendix
A.1 Difference in Human Behavior Experiments from the Existing Literature